### v1

# Exploring Graph Representations of Logical Forms for Language Modeling

Abstract
We make the case for language models over logical forms (LFLMs), arguing that such models are more data-efficient than their textual counterparts. To that end, we introduce the Graph-based Formal-Logical Distributional Semantics (GFoLDS) prototype, a pretrained LM over graph representations of logical forms, as a proof-of-concept of LFLMs. Using GFoLDS, we present strong experimental evidence that LFLMs can leverage the built-in, basic linguistic knowledge inherent in such models to immediately begin learning more complex patterns. On downstream tasks, we show that GFoLDS vastly outperforms textual, transformer LMs pretrained on similar amounts of data, indicating that LFLMs can learn with substantially less data than models over plain text. Furthermore, we show that the performance of this model is likely to scale with additional parameters and pretraining data, suggesting the viability of LFLMs in real-world applications.
Exploring Graph Representations of Logical Forms for Language Modeling
Michael Sullivan University at Buffalo Saarland University msullivan@lst.uni-saarland.de
1 Introduction
Although recent advances in LLMs have led to remarkable performance on a wide variety of benchmarks, the consistent improvements exhibited by SoTA LLMs are largely due to corresponding increases in model size Villalobos et al. (2022); Muennighoff et al. (2024). Given the Chinchilla Scaling Laws Hoffmann et al. (2022) and the rate at which SoTA LLMs are expanding, Villalobos et al. (2022) estimate that high-quality English training data will be exhausted at some point between 2026 and 2032: language model expansion is outpacing available natural language production. This suggests that—without models that use significantly less data than current approaches—LLMs’ performance increases will begin to decelerate substantially in the near future.
However, there is a considerable amount of evidence in the literature (e.g. Xu et al., 2021; Wu et al., 2021b; Prange et al., 2022; Sachan et al., 2021; Zhou et al., 2020; Zhang et al., 2020, 2022, etc.) indicating that linguistically-informed LMs—models whose inputs and/or architectures are augmented by linguistic knowledge—can improve performance without consuming more (text) data. In this paper, we argue for the use of LMs over logical forms (LFLMs): LMs that take as input semantic representations, rather than text. In particular, we posit the following hypothesis:
Figure 1: DMRS representation of the sentence “every bored person saw her yesterday.”
The Linguistic-Knowledge Catalysis Hypothesis (LKCH):
The (aspects of) linguistic knowledge incorporated into LFLMs obviates the need to learn elementary linguistic phenomena, allowing them to immediately begin learning more complex patterns.
One relevant corollary of the LKCH is that LFLMs can learn with less data: the basic linguistic knowledge built into LFLMs facilitates more rapid learning of advanced phenomena.
To make the case for LFLMs, we introduce the Graph-based Formal-Logical Distributional Semantics (GFoLDS; Section 3) model [^1] All code used in this paper available on GitHub: https://github.com/mjs227/GFoLDS as a proof-of-concept of such LMs. GFoLDS is a pretrained, encoder graph transformer Wu et al. (2021a) over structures derived from Dependency Minimal Recursion Semantics (DMRS; Copestake, 2009) representations (see Figure 1): directed, labeled acyclic graphs in which nodes correspond to predicates and edge labels denote relations (i.e. semantic roles) between them, in a similar manner to Abstract Meaning Representation (AMR; Banarescu et al., 2013) structures. Unlike AMR, which abstracts away from semantic features such as tense and number, DMRS includes these features, resulting in a more faithful representation of linguistic meaning.
The contributions of this work are threefold: firstly, we provide experimental support towards the validity of the LKCH, demonstrating that—from the start of pretraining—GFoLDS achieves near-peak performance on tasks designed to evaluate its elementary linguistic knowledge, and that this translates to more rapid learning of complex phenomena (Section 4).
Secondly, we demonstrate the viability of pretrained LFLMs in Section 5, by comparing the performance of GFoLDS to that of BERT (trained on $\sim$6.5 times more data than GFoLDS; Devlin et al., 2019) on a range of downstream tasks. Although the actual BERT models outperform GFoLDS, our model outperforms—by a wide margin—BERT models pretrained on the same data as GFoLDS, indicating that LFLMs can learn useful representations with much less data than their textual counterparts.
Thirdly, we establish that LFLMs have the potential to compete with textual LLMs at scale: in Section 6, we present evidence indicating that GFoLDS is likely to scale with respect to parameter count and pretraining dataset size.
2 Related Work
There exists a body of research indicating that injecting graph representations of linguistic structures into textual LMs can improve downstream performance. Xu et al. (2021) fuse dependency parse graphs into pretrainined transformer encoders (e.g. BERT and RoBERTa; Liu et al., 2019), and surpass the then-SoTA results on relation classification, entity typing, and question answering tasks, demonstrating the general utility of linguistically-informed LMs. Similarly, Wu et al. (2021b) inject syntactic dependency parse graphs into a pretrained BERT model, yielding then-SoTA results on semantic role labeling and relation extraction tasks. These authors additionally compare the respective impacts of syntactic and semantic representations on performance, and find that semantic representations are more beneficial with respect to downstream performance than syntactic structures.
Prange et al. (2022) show that linguistic structures can be incorporated into the input of GPT-2 (Radford et al., 2018) to improve next word prediction accuracy and entropy. Of particular relevance is their finding that Elementary Dependency Structures (EDS; Oepen and Lønning, 2006)—a semantic framework that is related to DMRS—yield greater performance improvements than syntactic (or other semantic) representations.
All of the models described thus far in this section are hybrid architectures that merge textual and graph representations. Furthermore, in these approaches, the textual component of the model is initialized from a pretrained LM such as BERT, RoBERTa, or GPT-2. This contrasts with the GFoLDS model, which takes only graph representations as input and is pretrained from scratch.
To the best of our knowledge, Functional Distributional Semantics at Scale (FDSAS; Lo et al., 2023) represents the only [^2] Aside from its predecessor: Functional Distributional Semantics Emerson (2018). extant LFLM aside from GFoLDS. FDSAS is a variational autoencoder over DMRS graphs that learns probabilistic truth-conditional functions for each predicate (node label). FDSAS is not, however, a transformer and so does not benefit from the flexibility afforded by such architectures: it is unclear how to scale the model to yield a deeper architecture, and it does not generate usable hidden states that can be passed to (for example) a classification head for downstream tasks. Additionally, its rigorous formal-semantic foundations give rise to a degree of inflexibility in this model: Lo et al. (2023) discarded prepositions, quantifiers, and modal verbs from FDSAS’ training dataset out of necessity.
3 GFoLDS
In this section, we describe the GFoLDS model architecture (Section 3.1), preprocessing steps that we employed on the DMRS graphs (Section 3.2), and the model’s pretraining procedure (Section 3.3).
3.1 Architecture
Figure 2: Top-level architecture of the GFoLDS model (top) and positional encoding network (bottom).
The GFoLDS model is a variant of the graph transformer paradigm Wu et al. (2021a), which was originally introduced for molecule graph classification. A graph transformer consists of a graph neural network (GNN) that encodes local neighborhood information, whose output is then fed to a permutation-invariant (i.e. without linear positional embeddings) transformer encoder for global message-passing (attention). Unique to this work is the GNN component of GFoLDS, which consists of an embedding layer and the positional encoding network (see Figure 2). The output of the embedding layer is fed into the positional encoding network, which provides each node with a representation of its local neighborhood in the DMRS graph structure.
Let $n_{i}$ denote the $i^{th}$ node in a graph $G$, and let $F(n_{i})$ be its set of DMRS features (person, number, tense, etc.; see Figure 7 in the Appendix). Then the output of the embedding layer $\vec{e_{i}}=E(X,G)_{i}$ (where $X$ denotes the node labels of $G$) is the sum of the embedding of the node’s label $\mathcal{E}_{T}(X_{i})$ with the normalized sum of the embeddings $\mathcal{E}_{F}(\phi)$ of each feature $\phi\in F(n_{i})$ (Equation 1).
$\vec{e_{i}}=\mathcal{E}_{T}(X_{i})+\text{{Norm}}\left(\sum_{\phi\in F(n_{i})}\mathcal{E}_{F}(\phi)\right)$ (1)
These summed feature and node/predicate embeddings are then passed to the positional encoding network (see Figure 2). This module consists of a linear layer (to project the embeddings from $d_{\text{{model}}}$ to $d_{\text{{SWA}}}$) followed by a stack of step-wise aggregation (SWA) layers (see Figure 3)—in which the output of each SWA layer is fed to the subsequent layer—followed by a second linear projection (to project from $d_{\text{{SWA}}}$ back to $d_{\text{{model}}}$).
Figure 3: Architecture of an SWA layer in the positional encoding network.
In their respective adaptations of the GraphSAGE (Hamilton et al., 2017) and Graph Convolutional Network (GCN; Kipf and Welling, 2017) architectures to directed graphs, Xu et al. (2018) and Tong et al. (2020) introduce forward and backward node projection layers, which encode information about incoming and outgoing connections (respectively) for a given node. In a similar fashion, each SWA layer contains a forward (Equation 2) and a backward SWA block (see Figure 3), which encode the nodes—and the semantic roles thereof—mapping into and out of a given node.
[…]
For each node $n_{i}\in G$, its forward representation $\vec{f_{i}}$ (i.e. the output of the forward SWA block) is the (normalized) sum of $W_{\ell}^{(f)}\vec{h_{k}}$ for each node $n_{k}$ with an edge $n_{k}\xrightarrow{\ell}n_{i}$ in the graph structure, where $W_{\ell}^{(f)}\in\mathbb{R}^{d_{\text{{SWA}}}\times d_{\text{{SWA}}}}$ is the forward edge projection linear layer for the edge label $\ell$. This representation of each edge label as a unique projection layer is conceptually similar to the label-specific matrices employed in Relational GCNs (Schlichtkrull et al., 2018) and Beck et al.’s (2018) adaptation of the Gated GNN (Li et al., 2016) architecture to labeled graphs.
The backward SWA block is architecturally identical to the forward block, but contains distinct edge projection layers $W_{\ell}^{(b)}$ and operates on the transpose $G^{T}$ of the graph. The respective forward and backward representations of each node $n_{i}$ are summed together and passed through a two layer, feed-forward block (identical to the feed-forward blocks in the model’s encoder layers).
The input to the encoder stack is then the sum of the outputs of the embedding layer and positional encoding network: $E(X,G)+P(E(X,G),G)$. This is analogous to (and inspired by) the approach taken by most transformer LMs (e.g. Devlin et al., 2019; Brown et al., 2020; Lewis et al., 2020; Raffel et al., 2020; Dubey et al., 2024, etc.), where the encoder stack input for token $t$ at position $p$ is the sum of the token embedding for $t$ and the positional embedding for $p$.
The encoder is not directly exposed to the graph structure: all nodes are able to attend to any other node(s). The encoder layers in the GFoLDS architecture are similar to those in BERT and Vaswani et al. (2017), with a few key differences with respect to the residual connections and layer normalization. A full description of the GFoLDS architecture is located in Appendix A.
3.2 Data Preprocessing
GFoLDS’ pretraining corpus consisted of $\sim$17.5 million randomly-selected sentences from the November 1, 2023 English Wikipedia dump [^3] https://huggingface.co/datasets/wikimedia/wikipedia; CC-BY-SA-3.0 license., constituting a total of $\sim$508 million words ($\sim$6.5 times smaller than BERT’s pre-training corpus). We first used Spacy’s SentenceRecognizer [^4] https://spacy.io/api/sentencerecognizer pipeline to extract individual sentences from the text. We then used the PyDelphin (Goodman, 2019) library with the ACE/ERG (Copestake and Flickinger, 2000) rule-based parser/grammar [^5] ERG-1214 release: https://github.com/delph-in/erg to obtain a DMRS representation of each sentence, before preprocessing the resulting DMRS structures to yield GFoLDS input graphs. The ACE/ERG parser was able to parse $\sim$84of the data, for a total of $\sim$14.6 million DMRS-derived graphs.
It was not feasible to tokenize named entities (CARGs) and out-of-vocabulary (OOV) items in the same manner as for the in-vocabulary DMRS predicates—i.e. by simply assigning an integer to each unique predicate string in the vocabulary. We therefore replaced all CARGS and OOV terms with the [MASK] token. These [MASK] tokens are not targets for prediction during the pre-training procedure (as they are OOV, so there is no possible target token); the goal is instead to have the model represent the OOV item with the closest in-vocabulary token, based on the context in which the OOV item appears. The removal of CARGs and OOV items is a stop-gap measure, and remains an open problem and barrier to the performance of the GFoLDS model: we defer the incorporation of CARGs and OOV items into the model’s input structures to future work (see Section 8.1).
Further details on the preprocessing procedures that we employed are located in Appendix B.
3.3 Pretraining
We pretrained GFoLDS with the masked-node modeling (MNM) objective, which is analogous to the MLM objective used to pretrain encoder transformer LMs. The model trained for four epochs, with a total training time of $\sim$102 hours ($\sim$25.5 hours per epoch) on a single NVIDIA A100 GPU. Further details on our pretraining procedure and hyperparameters are located in Appendix C.1.
We employed a GFoLDS model with two SWA layers and ten encoder layers (eight attention heads each), and set $d_{\text{{SWA}}}=d_{\text{{model}}}=1024$. The MNM prediction head that we used is identical to BERT’s MLM prediction head (aside from the difference in vocabulary size). This yields a total of $\sim$174 million parameters: for comparison, $\text{BERT}_{\text{base}}$ (12 encoder layers, $d_{\text{{model}}}$ = 768) and $\text{BERT}_{\text{large}}$ (24 encoder layers, $d_{\text{{model}}}$ = 1024) have $\sim$110 million and $\sim$335 million parameters, respectively.
4 Evaluating the LKCH
This section is dedicated to an investigation of the validity of the LKCH. The hypothesis can be broken down into two distinct claims: (i) that the (aspects of) linguistic knowledge incorporated into LFLMs obviates the need to learn elementary linguistic phenomena; and (ii) that this enables such models to then immediately begin learning more complex patterns.
We therefore divided this experiment into two parts, which respectively probe the model’s knowledge of elementary and complex linguistic phenomena. To validate the claim made in the LKCH that linguistically-informed LMs immediately begin to learn complex patterns, we evaluated the model at regular intervals throughout pretraining.
4.1 Comparison Models
As a baseline, we pretrained $\text{BERT}_{\text{base}}$ and $\text{BERT}_{\text{large}}$ (uncased) models from scratch on the same dataset as GFoLDS (the surface sentences), for the same number of epochs (four).
Recall that the ACE/ERG parser was only able to parse $\sim$84of the sentences in GFoLDS’ pretraining dataset (see Section 3.2). If GFoLDS were able to match the performance of BERT with $\sim$84of the pretraining data (for example), then—given the ACE/ERG parser’s $\sim$84successful parse rate—this model would provide no practical benefit over its textual counterparts. We therefore chose to pretrain the BERT comparison models (BERT-C) on GFoLDS’ entire pretraining dataset, rather than the parsable subset: the BERT-C models were pretrained with $\sim$1.19 times more data than GFoLDS.
We evaluated a variety of different pretraining hyperparameter configurations for BERT-C in order to yield the most rigorous comparison: a full description of these hyperparameter configurations is located in Appendix C.2.
4.2 Experimental Setup
This experiment consisted of two tasks that were designed to evaluate knowledge of elementary linguistic phenomena, and one to evaluate knowledge of more complex patterns. We evaluated GFoLDS and the comparison models on the three tasks at twenty evenly-spaced intervals per pretraining epoch, for a total of eighty points of comparison.
Assuming that the LKCH holds, we should expect to see GFoLDS outperform the BERT-C models on the complex task at each stage of the pre-training process, as—according to the hypothesis—GFoLDS is able to learn complex patterns faster than textual LMs. On the elementary tasks, we expect GFoLDS to outperform BERT-C, but also that GFoLDS’s performance will not improve significantly across pretraining, while BERT-C’s performance steadily increases: the LKCH asserts that an LFLM is already equipped with elementary linguistic knowledge, so its performance at the first checkpoint should be near (or above) that of a textual model such as BERT-C at the last checkpoint.
4.2.1 Complex Task
The complex task in this experiment is the RELPRON Rimell et al. (2016) dataset. This dataset consists of terms (nouns), each paired with a hypernym and up to ten properties: relative clauses that restrict that hypernym (see Table 2 in the Appendix). The development set consists of 65 terms and 518 properties ($\sim$8 properties per term on average), and the test set contains 73 terms and 569 properties ($\sim$7.8 per term). The task is to retrieve the properties that apply to each term, without including those that do not: the evaluation metric is Mean Average Precision (MAP) score. Due to the small size of the dataset (and the lack of a training split), the frozen, pre-trained models were used to obtain token probabilities for property ranking (see Appendix D for additional details).
To evaluate GFoLDS, we discarded all examples containing CARG-bearing predicates or OOV items (the “RELPRON-No-UNK/NE” row of Table 1). As discussed in Section 3.2, CARGs and OOV items are masked in the input graphs: given that each template only contains four content words (the hypernym, verb, relative clause subject/object, and the target term), the GFoLDS model is effectively blind to (at least) one third of the context in templates that contain OOV items or CARGs.
After discarding all examples that contain CARGs or OOV items, the resulting No-UNK/NE subset of the RELPRON test split contains 63 terms and 421 properties, for a total of $\sim$6.68 properties per term on average.
Figure 4: Precision scores on the two elementary tasks (left and center) and MAP scores on the RELPRON test set (right) across the 80 evenly-spaced training snapshots for GFoLDS and the BERT comparison models.
4.2.2 Elementary Tasks
POS-Prediction
The first task, POS-prediction, evaluates the LMs’ ability to model the distribution of parts-of-speech: this is a commonly-employed probing task used to assess LMs’ elementary linguistic competence (Waldis et al., 2024). We evaluated the models on 200 sentences (not drawn from their pretraining data) from English Wikipedia. For each sentence, we masked a single word and recorded its POS: we then extracted the models’ probability distribution over the masked word of each sentence, and recorded their precision at ten with respect to the masked word’s part-of-speech. For example, if the masked word was a noun, and nine of the model’s top ten most-likely predictions were nouns, then the model received a score of 0.9 for that instance.
Quantifier-Agreement
The second elementary task, quantifier-agreement, evaluates the models’ knowledge of quantifier number agreement—whether a quantifier restricts a singular or plural noun (or both)—across 179 sentences drawn from English Wikipedia (not used to pretrain the models). Although less common than POS prediction, quantifier agreement tasks have also been employed as a metric of language models’ elementary lingustic competence (see e.g. Huebner et al., 2021; Waldis et al., 2024).
As in the POS-prediction task above, we masked a single quantifier per sentence, then extracted the probability distribution over the masked quantifier of each sentence to compute precision at $k$. The value of $k$ varied as a function of the number of the noun in the masked quantifier’s restriction, as there were more quantifiers that restrict singular nouns ($k=14$) than plural nouns ($k=13$).
Additional details on the setup and evaluation of the elementary tasks are given in Appendix G.
4.3 Results
[…]
These results constitute strong evidence towards the LKCH. It is clear from this experiment that GFoLDS can model the elementary phenomena from the onset (and retains this ability throughout pretraining) and GFoLDS’ performance on the RELPRON test set suggests that this translates to more rapid learning of more complex patterns.
5 Downstream Tasks
In order to demonstrate the viability of LFLMs, we evaluated GFoLDS on a series of downstream tasks: our primary objective was to demonstrate that language models over logical forms are able to learn from less data than their textual counterparts. To that end, we compared GFoLDS to the BERT comparison (BERT-C) models introduced in Section 4.1, in order to show that GFoLDS is able to outperform a textual LM when pretrained on the same amount of data: it then follows that GFoLDS would be able to perform as well as a textual LM with less pretraining data.
We also compared GFoLDS to the original $\text{BERT}_{\text{base}}$ and $\text{BERT}_{\text{large}}$ (uncased) models, both of which are pretrained on roughly 6.5 times more data than GFoLDS.
5.1 Datasets
We evaluated the models on four downstream tasks: the RELPRON task described in Section 4, SNLI Bowman et al. (2015), the MegaVeridicality V2.1 factuality dataset White et al. (2018), and the McRae et al. (2005) property inference dataset.
5.1.1 SNLI
All five models were fine-tuned end-to-end on SNLI with a two-layer, feed-forward MLP classifier over their mean-pooled token/node embeddings. The details of our SNLI fine-tuning hyperparameters are located in Appendix E.
Graph Representations
[…]
Figure 5: Illustration of the derivation of $G(P_{i},H_{i})$ (bottom) from $G(P_{i})$ (top left) and $G(H_{i})$ (top right) for the contradiction example (the man is lying down, the man is standing).
RELPRON-All RELPRON-No-UNK/NE SNLI MegaVeridicality V2.1 McRae et al. (2005)
(MAP) (MAP) (Accuracy) (Accuracy) (Spearman $\rho$)
GFoLDS 0.651 81.0% 81.3% 0.205
$\text{BERT}_{\text{large}}$ 0.768 0.769 91.1% 85.6% 0.241
$\text{BERT}_{\text{base}}$ 0.667 0.690 90.7% 84.2% 0.247
[…]
FDSAS 0.580
Table 1: Results for GFoLDS and the four BERT models on downstream tasks.
[…]
5.1.2 MegaVeridicality V2.1
The objective in this task is to determine whether the event denoted by a subordinate clause is true, given the context of the matrix clause (see Table 3 in the Appendix). The MegaVeridicality V2.1 dataset consists of 5,026 examples, each with ten annotations of the factuality of the event: for this experiment, we converted the data into a binary classification task (see Appendix F).
As with RELPRON (see Section 4.2.1), we removed all examples containing OOV items and/or CARG-bearing predicates for this task, due to the lower amount of content words per example in MegaVeridicality V2.1 in comparison to SNLI. This left 3,126 remaining examples ($\sim$62%), of which we withheld 20% as a test set.
We fine-tuned all five models end-to-end for eight epochs with a two-layer (binary) classifier head over their mean-pooled embeddings, using a learn rate of $10^{-6}$, a weight decay value of $10^{-5}$, and a batch size of 16.
5.1.3 The McRae et al. (2005) Dataset
To evaluate the quality of GFoLDS’ embeddings in comparison to BERT, we evaluated the models on a property inference task using the McRae et al. (2005) feature norm database. This database consists of a set of 541 concepts (words) $W$ and 2,526 features $F$, in which each concept $w$ is assigned a feature vector $f(w)\in\mathbb{R}^{|F|}$: the value of $f(w)_{Q}$ is the value of the feature $Q$ for the word $w$ (see Table 4 in the Appendix).
Following Rosenfeld and Erk (2022), we created ten random folds consisting of 50 concepts each from the dataset. The concepts within each fold represent the set $U$ of unknown words, and those outside of the fold represent the set $K=W-U$ of known words. For each $u\in U$, its feature vector $f(u)$ is withheld: the task is to reconstruct $f(u)$.
As the goal of this experiment was to evaluate the quality of their embeddings, the models were not fine-tuned for this task: following Rosenfeld and Erk (2022), we used the Modified Adsorption Talukdar and Crammer (2009) graph label-propagation algorithm to estimate properties, using the cosine distance between two words’ embeddings to weight the graph edge between them.
The evaluation metric was Spearman’s $\rho$: we averaged over all scores for each unknown word in each fold to yield each model’s final score.
5.2 Results
The results of these experiments are given in Table 1. Although the two (original) BERT models outperform GFoLDS all four benchmarks, the two BERT comparison models both lag far behind our model, demonstrating that it is able to learn useful representations with less data than textual LMs across a wide range of downstream tasks.
[…]
6 Scalability
While the results of Section 5 show that GFoLDS is a viable competitor to textual models trained on similar amounts of data, this model is still outperformed by the original BERT models. It is therefore crucial to establish the scalability of GFoLDS: the degree to which we would expect its performance to scale if it were larger and/or pre-trained on more data. To that end, we applied the techniques of Muennighoff et al. (2024) to GFoLDS, to determine the degree to which GFoLDS is under- or over-parameterized—and therefore, by the scaling laws established in Muennighoff et al. (2024), over-/under-trained.
As loss is not an exact predictor of downstream performance (Shin et al., 2022; Tay et al., 2022; Xia et al., 2023), we follow Hoffmann et al. (2022) and evaluate the impact of pretraining token count on a validation task as well. We again used the RELPRON dataset for this purpose, as SNLI introduces potential confounding factors in the form of the fine-tuning procedure.
6.1 Experimental Setup
We pretrained five GFoLDS models on 50%, 25%, 12.5%, 6.25%, and 3.125% of the pretraining data used in Section 3.3. Following Muennighoff et al. (2024), we ensured that the iterations pretrained on less data always use a randomly-selected subset of the dataset used in those with more data. Aside from the differing number of pretraining tokens, all models were pretrained using the same procedure and hyperparameters as described in Section 3.3, as the focus of this experiment was the effect of training tokens with a fixed parameter count
6.2 Results
Figure 6 shows that final pretraining loss consistently decreases with the number of pretraining tokens from 3.125% to 50% of the data. After this point, the final loss value plateaus: the 50% run (1.3232) finishes with a lower cross-entropy loss than the actual (100%) run (1.3331), although this slight difference is likely explained by the noise introduced by random initialization of the models’ parameters. In Appendix H, we prove that—assuming that an analogue of the Muennighoff et al. (2024) scaling laws holds for GFoLDS—it can only be the case that the final loss for the 100% run is (roughly) equal to that for the 50% run if GFoLDS is underparameterized for both the 100% run and the 50% run. This is to say that GFoLDS requires much less pretraining data per parameter than textual models: the Chinchilla Scaling Laws Hoffmann et al. (2022) predict that a textual LLM with the same parameter count ($\sim$174 million) as GFoLDS necessitates $\sim$5 billion pretraining tokens—roughly twenty times more than the $\sim$254 million tokens at which GFoLDS becomes overparameterized.
Figure 6: Final pretraining loss (left) and RELPRON MAP (right; No-UNK/NE test split). The 100% run denotes GFoLDS pretrained on the entire dataset.
We also observe improvement of +0.082 in RELPRON MAP score from the 50% (0.569) to the 100% (0.651) run—higher than the +0.062 increase from the 25% (0.507) to the 50% run.
The results of this experiment lead to the conclusion that GFoLDS is likely scalable in terms of model size and pretraining data. The application of the Muennighoff et al. (2024) scaling laws to GFoLDS indicates that increasing the parameter count will decrease final pretraining loss, while its run-over-run performance on the RELPRON test set suggests that larger pretraining datasets will lead to corresponding increases in the model’s downstream performance.
7 Conclusion
In Section 4, we provided direct evidence in support of the LKCH: critically, this finding indicates that LFLMs can learn with less data than textual models. This is supported by our findings in Section 6, which indicate GFoLDS is underparameterized with half of its pretraining data.
To the best of our knowledge, the experiments in Section 5 represent the first time that a language model pretrained solely over logical forms has been evaluated on a wide range of downstream tasks. GFoLDS’ capacity to be applied to tasks ranging from RELPRON to SNLI (see Section 5) demonstrates this model’s versatility, and these results therefore represent a significant step towards demonstrating the practical viability of LFLMs.
Although much work remains to be done before LFLMs can reach the same level of performance and utility as their textual counterparts (see Section 8), the results of Sections 4-6 suggest that such models present an exciting avenue for continuing the improvement of language models at a more sustainable rate of data consumption than LLMs over text.
8 Limitations
This section is divided into two parts: Section 8.1 overviews the limitations of the current GFoLDS pipeline laid out in Section 3, while Section 8.2 discusses the limitations of the experiments conducted in Sections 4-6.
8.1 Limitations of the Current Approach
Tokenization
As discussed in Section 3.2, we replaced all OOV node labels with the [MASK] token. Although this preprocessing step was necessary due to the tokenization scheme that we employed, it prevents the model from seeing every word in every input sentence. For the same reason, we removed CARGs (named entities), keeping only the CARG-bearing predicates (nodes): this amounts to, for example, replacing the sentence “John went to the park in the spring of 2017” with “[NAMED] went to the park in the [SEASON] of [YEAR].”
The removal of these items from its DMRS-derived input graphs almost certainly negatively impacts GFoLDS’ performance on downstream tasks, and entirely precluded a direct comparison with FDSAS on the RELPRON benchmark (see Section 5.2). However, this limitation lies purely with the current GFoLDS architecture—rather than the proposed approach in and of itself—as we simply tokenized the nodes by assigning a unique integer to each in-vocabulary term (i.e. node label). The ACE/ERG parser can still parse OOV items: the node label of an OOV term is constructed by appending the suffix “_unknown” (along with a part-of-speech tag) to the corresponding string in the surface text.
Future work in this area must involve redesigning the model’s approach to node embeddings in order to overcome this limitation, for example by replacing GFoLDS’ embedding layer with a small, character-level encoder transformer that is applied to each individual node. As we intended this work to primarily be a proof-of-concept of LFLMs, we left this issue to future research in order to retain focus on that goal.
Multiple-Sentence Inputs
For the SNLI dataset, we were able to overcome GFoLDS’ inability to process sequences of multiple sentences by converting each premise/hypothesis pair into a single, connected graph. However, this task-specific remedy cannot be readily generalized to other multiple-sentence NLP benchmarks. In future work, we intend to address this deficiency through sentence-level positional encodings (e.g. rotary embeddings; Su et al., 2024): a sequence of sentences represented as the disjoint union of their DMRS graphs, where the positional encoding $\vec{p_{i}}$ is added (or concatenated) to the embedding of each node in the $i^{th}$ graph in the sequence.
DMRS Parsing
Although the ACE/ERG parser that we employed to derive DMRS graphs from natural language enjoys high precision (93.77%; Zamaraeva and Gómez-Rodríguez, 2024), it suffers from relatively low recall: only $\sim$84of our pretraining dataset was able to be parsed. This can (and should) be addressed in future work, perhaps through the use of a neural DMRS parser (e.g. Buys and Blunsom, 2017), which trades lower precision for higher recall than a rule-based parser such as ACE/ERG.
Additionally, the grammar of the ACE/ERG parser (ERG: English Resource Grammar) is solely an English grammar, rendering our approach inapplicable to other languages [^7] Although broad-coverage grammars do exist for some other higher-resource languages (e.g. the Spanish Resource Grammar; Zamaraeva et al., 2024).
Model Type
The GFoLDS model described in this work is an encoder model: in order to be able to perform the same range of tasks as current SoTA LLMs, LFLMs must have generative capabilities. In future research, we intend to adapt GFoLDS to construct a sentence-level, graph-to-graph generative LFLM, by adapting existing work in the domain of molecule graphs on autoregressive graph generation (e.g. Bacciu et al., 2020; Goyal et al., 2020; Bacciu and Podda, 2021) that we believe can be readily extended to DMRS, using (for example) a breadth-first search to impose a canonical generation order on the graph nodes during training.
8.2 Experimental Limitations
Logical Representations
While the GFoLDS model demonstrates that language models over DMRS representations are able to learn with less data than textual models, we did not evaluate the LKCH using other logical-form representations such as AMR. AMR in particular is a problematic representational format for the GFoLDS model, as this framework utilizes $\sim$100 edge labels Banarescu et al. (2013) (our DMRS-derived graphs have nine; see Table 5 in the Appendix). Recall from the discussion in Section 3.1 that each unique edge label corresponds to unique forward/backward edge projection layers within each SWA layer (see also Appendix B): with the architectural configuration of the GFoLDS model used in this work (i.e. two SWA layers with $d_{\text{{SWA}}}=1024$), the additional AMR edge labels would add $\sim$380 million parameters to the positional encoding module (the model in this work has only 174 million total parameters).
Although they may require significant architectural modifications, future work should investigate the use of AMR, EDS, and other graph-based semantic representations for language modeling, in order to determine the most effective framework for this application.
Dataset Size
Due to computational limitations, we did not investigate the GFoLDS model at scale (i.e. with the same amount of training data as BERT): it required roughly one month to parse the training dataset used in this work into DMRS [^8] We do not view this as a severe limitation of the proposed approach: DMRS parsing can be achieved in parallel (each parse is independent of the others), and so the speed of parsing scales linearly with the number of CPU cores. With sixteen times as many cores as used in this work, for example, a dataset the size of BERT’s pretraining data could be parsed in roughly two weeks.. Although this does introduce the risk that our proposed method may only show superiority on a small-scale dataset, the results of Section 6 indicate that this risk is relatively low: extrapolating from the pattern shown in Figure 6 (see the discussion in Section 6.2), we would expect GFoLDS to surpass $\text{BERT}_{\text{large}}$ (the original model) on the RELPRON test set with four times the current training data (still less than that of $\text{BERT}_{\text{large}}$, which additionally has twice the parameter count).
Model Size
We did not investigate the effect of parameter size on the GFoLDS model: these experiments were only performed with a 174 million parameter LFLM.
Appendix A Encoder Architectural Details
As discussed in Section 3.1, the outputs of the embedding layer and positional encoding network ($E(X,G)$ and $P(E(X,G),G)$, respectively) are summed together and passed to the encoder stack.
The encoder layers in the GFoLDS architecture are similar to those in BERT and Vaswani et al. (2017) [^9] Which is also the architecture employed in Wu et al.’s (2021a) original formulation of the graph transformer., but contain a few key differences—in particular with respect to the residual connections and interrelated layer normalization. As shown in Figure 8, in a BERT encoder layer, the layer input is directly passed to the multi-head attention module, immediately followed by a skip connection and layer normalization. The output of this first layer normalization is then passed to the feed-forward layer, which is again immediately followed by a skip connection and second layer normalization.
Figure 8: Architecture of a GFoLDS (top) and BERT (bottom) encoder layer.
In a GFoLDS encoder layer, the input is first layer-normalized before being passed to the multi-head attention module, which is followed by a skip connection. The next skip connection—in contrast to BERT—is copied before layer normalization, which itself is followed by the feed-forward layer (which is identical to a BERT encoder feed-forward layer) and a second skip connection; this skip connection is not followed by layer normalization.
These architectural differences are motivated by the fact that—since the introduction of BERT—normalization outside of the residual connection (i.e. $\text{{Norm}}(x+f(x))$) has been shown to be problematic. Godbole et al. (2023) instead recommend normalization inside the residual (i.e. $x+f(\text{{Norm}}(x))$), which we implemented in GFoLDS.
Appendix B Data Preprocessing
Original Label Interpretation/Role Replacement
ARG1 First-place argument
ARG2 Second-place argument
ARG3 Third-place argument
ARG4 Fourth-place argument
MOD Indicates a shared handle
between two predicates
RSTR Restriction of a quantifier
ARG Argument of the “unknown MOD
predicate
L-INDEX Left-hand conjunct of two INDEX
coordinated variables
R-INDEX Right-hand conjunct of two INDEX
coordinated variables
L-HNDL Left-hand conjunct of two HNDL
coordinated handles
R-HNDL Right-hand conjunct of two HNDL
coordinated handles
Table 5: DMRS edge labels (left), the role that they play in describing meaning (center), and the edge labels that they are replaced with (right) during preprocessing (if any: “ — ” indicates that a label is retained after preprocessing).
After removing CARGs and OOV items from the graph structures (see Section 3.2), there remain a few—relatively less significant—preprocessing steps that we took in order to transform DMRS representations into inputs for the GFoLDS model.
Figure 9: DMRS representation of the noun phrase “the large building”.
[…]
While this representational choice is sensible from the perspective of formal semantics, it is undesirable from the viewpoint of machine learning. The fact that ARG only links unknown to other predicates (and is always included when unknown is in the graph) makes that predicate extremely predictable: if, during pretraining, the model is given a graph with a masked unknown predicate, it needs only look for the ARG-labeled edge to know that unknown is the masked node. This means that the model does not need to learn any co-occurrence relations between unknown and other nodes in the graph in order to learn to reliably predict the distribution of unknown.
Second, recall that each unique argument label is assigned unique forward and backward $d_{\text{{SWA}}}\times d_{\text{{SWA}}}$ edge projection layers in each SWA layer (see Section 3.1). This is to say that each unique edge label corresponds to $2n(d_{\text{{SWA}}})^{2}$ parameters in the GFoLDS model, where $n$ denotes the number of SWA layers. If, for example, $n=2$ and $d_{\text{{SWA}}}=1024$, then each additional edge label adds 4,194,304 parameters to the model architecture. Given the highly specialized function of the ARG edge label, it seems rather unreasonable to allocate so many parameters to its representation.
Therefore, during preprocessing, we converted each ARG label to the MOD label (see Table 5): another purely structural DMRS edge label that is used to indicate handle equality between predicates when other argument-label edges alone are insufficient to do so (Muszynska, 2020).
[…]
Finally, we removed from the graph structures all instances of focus_d and parg_d: predicates with a purely discourse-pragmatic role, which indicate that the predicates that they modify are focus-topicalized or the subject of a passivized verb (respectively).
Appendix C Pretraining Details
This section describes additional details and hyperparameters of the pretraining processes for GFoLDS (Section C.1) and the BERT comparison models used in Sections 5-4 (Section C.2).
C.1 GFoLDS
Pretraining Objective
During BERT’s pretraining procedure, while 15% of the input tokens are selected for prediction, only 80% of the selected tokens are masked: this is to account for the mismatch between the model’s pretraining and fine-tuning distributions that arises from the fact that the [MASK] token only occurs during pretraining. However, for GFoLDS, the [MASK] token does occur during fine-tuning as well, due to OOV items (as discussed in Section 3.2). Furthermore, Wettig et al. (2023) find that higher selection rates—and higher masking rates—result in improved performance on downstream tasks, when compared to the selection and masking/replacement rates reported in Devlin et al. (2019). For these reasons, we chose to mask 100% of the selected tokens during pre-training, with the slightly higher selection probability of 20%.
The MNM prediction head that we used is identical to BERT’s MLM prediction head (aside from the difference in vocabulary size): a $d_{\text{{model}}}\times d_{\text{{model}}}$ linear layer, followed by GeLU activation, layer norm, and a $d_{\text{{model}}}\times\text{22077}$ (the size of the vocabulary) linear layer.
Hyperparameters
We pretrained GFoLDS with a batch size of 16 for four epochs with the AdamW optimizer Loshchilov and Hutter (2017) and a weight decay value of $10^{-5}$. We set an initial learn rate of $10^{-5}$, with a linearly interpolated learn rate between values of $2\times 10^{-5}$ at the end of the first epoch, $10^{-5}$ at the end of the second, $3\times 10^{-6}$ at the end of the third, and $10^{-6}$ at the end of the fourth. That is to say that the learn rate increased linearly from $10^{-5}$ to $2\times 10^{-5}$ during the first epoch, decreased linearly from $2\times 10^{-5}$ to $10^{-5}$ during the second epoch, and so on.
Figure 10: GFoLDS pretraining cross-entropy loss.
At the specified batch size of 16, the model trained at a rate of roughly 25 hours and 36 minutes per epoch on a single NVIDIA A100 GPU, for a total training time of 102 hours and 24 minutes. GFoLDS converged to a cross-entropy loss of $\sim$1.3331 at the end of the fourth epoch (see Figure 10).
C.2 BERT Comparison Models
Figure 11: Learning rates (left), cross-entropy loss values (center), and MAP scores on the RELPRON development split (right) across training steps for the $\text{BERT}_{\text{base}}$ pretraining trials (a)-(c).
Given the differences in size and modality between the BERT and GFoLDS models, the best-performing set of pretraining hyperparameters for BERT on this dataset is not likely to be identical to those of GFoLDS. We therefore evaluated a variety of different hyperparameter configurations for BERT in order to yield the most equitable comparison with GFoLDS. Due to the relatively higher cost associated with training $\text{BERT}_{\text{large}}$ (over three times larger than $\text{BERT}_{\text{base}}$), we performed the majority of the trials with $\text{BERT}_{\text{base}}$, then transferred the best-performing configuration found during these experiments to $\text{BERT}_{\text{large}}$.
We first evaluated $\text{BERT}_{\text{base}}$ on three different configurations across which the learn rate schedule, weight decay, and masking rates varied: all three configurations employed the next sentence prediction (NSP) secondary pretraining task Devlin et al. (2019). Due to hardware constraints, we were limited to using a batch size of 16 for all of the BERT pretraining trials.
The first configuration (a) was identical to that which we employed for GFoLDS (see Appendix C.1): a weight decay value of $10^{-5}$; a linearly interpolated learn rate between values of $2\times 10^{-5}$ at the end of the first epoch, $10^{-5}$ at the end of the second, $3\times 10^{-6}$ at the end of the third, and $10^{-6}$ at the end of the fourth (with an initial learn rate of $10^{-5}$); and a token selection probability of 20% with a masking probability of 100%.
In the second configuration (b), we used a hyperparameter configuration that was identical to that of the original BERT models: a weight decay value of $10^{-2}$; linear learn rate warmup to $10^{-4}$ across the first 1% of the training run (with linear decay thereafter); and a token selection probability of 15% with a masking and replacement rates of 80% and 10%, respectively.
However, the original BERT models were pretrained with a batch size of 256, while trials (a) and (b) use a batch size of 16. Although we were not able to increase the batch size due to hardware constraints (as mentioned above), Granziol et al. (2022) show that proportional (to batch size) learn rate scaling can be used to control for the effect of batch size on training loss. We therefore introduced a third trial (c): this configuration was identical to that of (b) with the exception of the peak learn rate value, which we scaled down to $2\times 10^{-5}$ to account for the difference in batch size.
We then evaluated trials (a)-(c) with respect to learn rate and a validation task that does not require fine-tuning: the development split of the RELPRON (Rimell et al., 2016) dataset (see Section 4.2.1). The results of these experiments are shown in Figure 11.
The model clearly failed to learn with the original BERT pretraining hyperparameters (b)—likely due to the mismatch in batch size discussed above—and finished training with a minimum cross-entropy loss of 7.8864 and a peak MAP score of 0.040 on the RELPRON development set. Of the two remaining configurations, (c) outperforms (a) both in terms of minimum cross-entropy (4.2988 vs. 4.8716) and peak MAP score (0.207 vs. 0.137).
Liu et al. (2019) suggest that pretraining with the secondary NSP objective does not improve (and in some cases may even hinder) model performance. We therefore conducted a fourth hyperparameter trial with $\text{BERT}_{\text{base}}$, using the same configuration as in (c) above, but excluding the NSP task. The results of this experiment are shown in Figure 12.
Figure 12: Cross-entropy loss values (left), and MAP scores on the RELPRON development split (right) across training steps for the $\text{BERT}_{\text{base}}$ pretraining configuration (c) with and without the secondary NSP objective.
The variant of configuration (c) without NSP outperforms the original trial in terms of cross-entropy loss (3.8339 vs. 4.2988, respectively; see Figure 12)—this is to be expected: the loss values reported for the variant with NSP are the sum of the NSP loss with the masked language modeling (MLM) loss. However, the non-NSP configuration also outperforms its NSP counterpart in terms of peak MAP score on the RELPRON development set: 0.267 vs. 0.207 (respectively). For this reason, we selected the non-NSP variant of the model pretrained with hyperparameter configuration (c) as the $\text{BERT}_{\text{base}}$ comparison model to be used in the experiments in this work.
We then pretrained $\text{BERT}_{\text{large}}$ with (non-NSP) hyperparameter configuration (c). As shown in Figure 13 (configuration large-(a)), this model failed to converge: it finished training with a minimum cross-entropy loss of 7.1856 and a peak MAP score of 0.039 on the RELPRON development set. As larger neural networks are more prone to overfitting (Caruana et al., 2000; Salman and Liu, 2019), we scaled the peak learn rate by a factor of $1/10$ in trial large-(b) to account for the difference in size between the $\text{BERT}_{\text{base}}$ and $\text{BERT}_{\text{large}}$ models.
Figure 13: Learning rates (left), cross-entropy loss values (center), and MAP scores on the RELPRON development split (right) across training steps for the $\text{BERT}_{\text{large}}$ pretraining trials (a)-(b).
Trial large-(b) vastly outperformed large-(a) in terms of both minimum cross-entropy loss (5.1998 vs. 7.1856) and peak RELPRON development split MAP score (0.039 vs. 0.083), although it trails far behind the best-performing $\text{BERT}_{\text{base}}$ configuration by both metrics (3.8339 vs. 5.1998 cross-entropy; 0.267 vs. 0.083 RELPRON MAP). This substantial difference in performance between the base and large BERT variants is to be expected: it is well-known that larger neural networks require more training data in order to properly converge (see e.g. Hoffmann et al., 2022; Muennighoff et al., 2024), and $\text{BERT}_{\text{large}}$ has over triple the amount of parameters (and double the number of encoder layers) as the base version of the model. Further compounding this issue is the fact that we pretrained these comparison models on 6.5 times less data (and for ten times fewer epochs) than was intended for the BERT models.
With a batch size of 16, $\text{BERT}_{\text{large}}$ trained at a rate of roughly 35 hours and 20 minutes per epoch on a single NVIDIA H100 GPU, for a total training time of 141 hours and 20 minutes. For comparision, GFoLDS and $\text{BERT}_{\text{base}}$—each trained on an NVIDIA A100—required 102 hours and 24 minutes, and 46 hours and 36 minutes, respectively (total training time).
Appendix D RELPRON Task Setup
To evaluate the models on RELPRON, we constructed templates out of each (term, hypernym, property) triple: for example, the triple (telescope, device, astronomers use) yields the template “a device that astronomers use is a telescope” (see Table 2). We then replaced the target term with the [MASK] token (e.g. “a device that astronomers use is a [MASK]”): the probability assigned to a given term under the masked distribution is taken as proportional to the probability that the property applies to that term.
Appendix E SNLI Fine-Tuning Details
We fine-tuned GFoLDS and $\text{BERT}_{\text{base}}$ for five epochs with a batch size of 16, a weight decay value of $10^{-5}$, an initial learn rate of $10^{-5}$, and a linearly-interpolated learn rate (updated at each batch) between values of $2\times 10^{-5}$ at the end of the first epoch, $3\times 10^{-5}$ at the end of the third, $10^{-6}$ at the end of the fourth, and $10^{-7}$ at the end of the fifth. We fine-tuned $\text{BERT}_{\text{large}}$ with identical hyperparameters, except all learn rates mentioned above were multiplied by 1/10 for this model ($\text{BERT}_{\text{large}}$ was unstable with the higher learn rate used for GFoLDS and $\text{BERT}_{\text{base}}$).
The fine-tuning hyperparameters for the BERT models were admittedly not selected using as rigorous of a search procedure as that employed during pretraining (see Appendix C.2). However, the accuracy that the (original) $\text{BERT}_{\text{base}}$ and $\text{BERT}_{\text{large}}$ models attained with the hyperparameter configurations described in the above paragraph (see Table 1) matched that reported for those models on the SNLI dataset in Zhang et al. (2020). It is therefore reasonably safe to assume that this set of hyperparameters is (near-)optimal for the BERT models with respect to this data.
Type Quantifiers
SG another, either, neither, that, this, every, a(n), each
PL these, certain, most, those, all, such, both
BOTH some, the, any, enough, no, which
Table 6: Quantifier types—along with a list of the quantifiers belonging to each type—used in the quantifier-agreement task.
Appendix F MegaVeridicality V2.1 Task Setup
We converted this dataset to a binary classification task by assigning a values of $1$, 00, and $-1$ to the labels yes, maybe, and no (respectively). We then assigned each example a value of 1 (i.e. the subordinate event is portrayed as true) if its mean value was greater than zero, and 0 otherwise.
Appendix G Elementary Tasks
POS-prediction
As discussed in Section 4.2, we evaluated the models on 200 sentences drawn from English Wikipedia. We first parsed each sentence using the ACE/ERG DMRS parser (see Section 3.2), which automatically labels the part-of-speech of each predicate in the DMRS representation of a sentence. We then randomly selected a single word to mask from each parsed sentence—subject to the condition that the selected word must be mapped to a single token by the BERT tokenizer (in order to facilitate the evaluation of the BERT comparison models)—and recorded to part-of-speech of the selected word. This resulted in a dataset consisting of 58 masked-quantifier, 28 masked-preposition, 33 masked-verb, 63 masked-noun, and 18 masked-adjective sentences.
Determining the part-of-speech for each prediction of the GFoLDS model was trivial: DMRS predicates include part-of-speech tags, so we simply checked the tag of each predicted predicate. The BERT models, however, necessitated the use of the NLTK POS tagger [^11] https://www.nltk.org/api/nltk.tag.pos_tag.html. For each sentence $s$, and each of the model’s top-ten predicted tokens $w$ for $s$, we created a new sentence $s_{w}$ by replacing the masked word of $s$ with the prediction $w$, and ran the NLTK POS tagger over $s_{w}$ to obtain the tag for $w$. Note that, while the NTLK POS tagger is not perfect, it does achieve 95+accuracy on English-language data (Jacobsen et al., 2021), and therefore is sufficiently robust to yield an estimate of the model’s performance.
We chose to use (bounded) precision as the evaluation metric for this task because of the large amounts of positive examples for each class (especially nouns, verbs, and adjectives), which precluded the calculation of metrics that incorporate false negatives (e.g. recall and F1). We recorded each model’s mean precision across all 200 sentences as its final score for this task.
Quantifier-agreement
We first parsed each sentence with the ACE/ERG parser, which explicitly labels the number of each noun: this allowed the automatic extraction of the number of the noun in the restriction of a given quantifier. We then randomly selected a single quantifier from each sentence to mask, and recorded the number of the noun in the quantifier’s restriction.
We sorted all of the quantifiers into one of three categories/types (see Table 6): singular (can only restrict singular nouns), plural (can only restrict plural nouns), and both (can restrict either kind of noun). Note that the both-type quantifiers were used only for evaluation: the type of the masked quantifiers was recorded only as singular or plural—all nouns are either singular or plural, and the target type was determined by the number assigned by the ACE/ERG parser to the noun in the quantifier’s restriction.
When computing precision for this task, all non-quantifier words in the models’ top-$k$ predictions were treated as false positives, while both-type quantifiers in the top $k$ were treated as true positives, regardless of the target type (singular or plural). As in the POS-prediction task above, we recorded each model’s mean precision across all 179 sentences as its final score for this task.
Appendix H Scalability
H.1 Background and Notation
Hoffmann et al. (2022) investigate the relationship between model size (i.e. number of parameters) and amount of pretraining data, and the final pretraining cross-entropy loss of LLMs. The optimal number of parameters and amount of data (number of tokens) for a fixed compute budget—expressed in terms of floating-point operations (FLOPs)—is given in Equation 3, where $N_{opt}$ denotes the optimal number of model parameters, $D_{opt}$ the optimal number of training tokens, $C$ the compute budget, and $L(N,D)$ the model’s final pretraining loss on $D$ tokens with $N$ parameters.
[…]
$L(N,D)\approx\hat{L}(N,D)=E+\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}$ (4)
However, Hoffmann et al. (2022) only consider unique training data—i.e.  pretraining a language model for a single epoch. Muennighoff et al. (2024) extend Hoffmann et al.’s (2022) experiments to the case of repeated training data: pretraining the model for multiple epochs on the same dataset. The authors fit an analogous loss-prediction function to that given in Equation 4 (Equation 5a), $\hat{L}^{(r)}(N,U_{D})$, which estimates the final loss for a language model with $N$ parameters trained on $U_{D}$ tokens for $r$ repetitions (i.e. epochs).
$L^{(r)}(N,U_{D})\approx\hat{L}^{(r)}(N,U_{D})=E+\frac{A}{\hat{N}^{\alpha}}+\frac{B}{\hat{D}^{\beta}}$ (5a)
$\hat{D}=U_{D}+U_{D}R^{*}_{D}(1-e^{-r/R^{*}_{D}})$ (5b)
$\hat{N}=U_{N}+U_{N}R^{*}_{N}(1-e^{-R_{N}/R^{*}_{N}})$ (5c)
$U_{N}=\text{{min}}\{N,N_{\text{{opt}}}(U_{D})\}$ (5d)
$R_{N}=\frac{N}{U_{N}}-1$ (5e)
[…]
H.2 Proof
The term $D$ in Equation 5 denotes the number of pretraining tokens. This metric is likely not a perfect predictor for GFoLDS: recall from Section 3.1 that the model contains separate projection layers for each DMRS edge label type. The model therefore receives training signal not only from the node labels (i.e. tokens), but also from the edge labels. For example, an input graph with six tokens and five edges will update fewer model parameters than a graph with six tokens and ten edges.
It is beyond the scope of this work to establish exact scaling laws for GFoLDS and determine the graph-based analogue to the term $U_{D}$ in Equation 5. However, it is clear that such a $U_{D}$ value scales (more or less) linearly with the number of graphs in the pretraining dataset: as with a textual model, $U_{D}$ can be expressed as the sum of the amounts of data (regardless of how it is quantified) contributed by each individual input graph (respectively, sequence) in the dataset.
It is therefore reasonable to assume that $U_{D^{\prime}}\approx U_{D}/2$, where $U_{D}$ denotes the dataset for the 100% run, and $U_{D^{\prime}}$ that of the 50% run (see Section 6). We may express the relationship between the 50% run loss and the 100% loss in the notation introduced in Appendix H.1 (while leaving the exact definition of $U_{D}$ to future work): $L^{(4)}(N,U_{D})\approx L^{(4)}(N,U_{D}/2)$. Recall from Equation 5a that the term $E$ is a constant, and so can be ignored. We are then left with the following (approximate) equality in Equation 6a.
$\frac{A}{\hat{N}^{\alpha}_{1}}+\frac{B}{\hat{D}^{\beta}_{1}}\approx\frac{A}{\hat{N}^{\alpha}_{2}}+\frac{B}{\hat{D}^{\beta}_{2}}$ (6a)
$\hat{D}_{1}=U_{D}+U_{D}R^{*}_{D}(1-e^{-4/R^{*}_{D}})$ (6b)
$\hat{D}_{2}\approx(U_{D}/2)+(U_{D}/2)R^{*}_{D}(1-e^{-4/R^{*}_{D}})$ (6c)
$\hat{N}_{1}=U^{(1)}_{N}+U^{(1)}_{N}R^{*}_{N}(1-e^{-R^{(1)}_{N}/R^{*}_{N}})$ (6d)
$\hat{N}_{2}=U^{(2)}_{N}+U^{(2)}_{N}R^{*}_{N}(1-e^{-R^{(2)}_{N}/R^{*}_{N}})$ (6e)
$U^{(1)}_{N}=\text{{min}}\{N,N_{\text{{opt}}}(U_{D})\}$ (6f)
$U^{(2)}_{N}\approx\text{{min}}\{N,N_{\text{{opt}}}(U_{D}/2)\}$ (6g)
$R^{(k)}_{N}=\frac{N}{U^{(k)}_{N}}-1$ (6h)
Where the left-hand expression in Equation 6a corresponds to the 100% run ($\hat{L}^{(4)}(N,U_{D})$), and the right-hand expression corresponds to the 50% run ($\hat{L}^{(4)}(N,U_{D}/2)$). Note that $\hat{D}_{2}\approx\hat{D}_{1}/2$ (Equation 7).
$\begin{split}\hat{D}_{2}&\approx\frac{U_{D}}{2}+\frac{U_{D}}{2}R^{*}_{D}(1-e^{-4/R^{*}_{D}})\\ &=\frac{U_{D}+U_{D}R^{*}_{D}(1-e^{-4/R^{*}_{D}})}{2}=\frac{\hat{D}_{1}}{2}\end{split}$ (7)
Given the vast architectural and modal differences between GFoLDS and textual transformer models, we do not assume that the coefficients $E$, $A$, $B$, $\alpha$, $\beta$, $R^{*}_{D}$, and $R^{*}_{N}$ fitted in Muennighoff et al. (2024) are identical for GFoLDS. Moreover, as stated above, it is beyond the scope of this work to establish exact scaling laws for this model. However, from the fact that final loss decreases as pretraining data increases from the 3.125% to the 50% runs (see Section 6.2), we know that the coefficients $B$ and $\beta$ in 6 must be positive: this—in conjunction with the (approximate) equality in Equation 7—means that it cannot be the case that $B/\hat{D}^{\beta}_{1}>B/\hat{D}^{\beta}_{2}$.
Assume that $B/\hat{D}^{\beta}_{1}\ll B/\hat{D}^{\beta}_{2}$: given the equality in Equation 6a, it must then be the case that $A/\hat{N}^{\alpha}_{1}-A/\hat{N}^{\alpha}_{2}\approx B/\hat{D}^{\beta}_{2}-B/\hat{D}^{\beta}_{1}$, and therefore that $A/\hat{N}^{\alpha}_{2}\ll A/\hat{N}^{\alpha}_{1}$.
Assume further that the model is underparameterized at 100% of the data (i.e. that $N<N_{\text{{opt}}}(U_{D})$): then $U^{(1)}_{N}=\text{{min}}\{N,N_{\text{{opt}}}(U_{D})\}=N$ and $R^{(1)}_{N}=(N/U^{(1)}_{N})-1=0$, which implies that $\hat{N}_{1}=U^{(1)}_{N}+U^{(1)}_{N}(1-1)=U^{(1)}_{N}=N$ (see Equation 6d). As the terms $A$ and $\alpha$ are constants in Equation 5a (and therefore in Equation 6a), we may reduce to the inequality expressed in Equation 8 (where the last logical equivalence is by definition of $R^{(2)}_{N}$; see Equation 6h).
$\begin{split}&\hat{N}^{-\alpha}_{1}\gg\hat{N}^{-\alpha}_{2}\leftrightarrow\hat{N}_{1}\ll\hat{N}_{2}\\ &\leftrightarrow N\ll U^{(2)}_{N}+U^{(2)}_{N}R^{*}_{N}(1-e^{-R^{(2)}_{N}/R^{*}_{N}})\\ &\leftrightarrow\frac{N-U^{(2)}_{N}}{U^{(2)}_{N}}\ll R^{*}_{N}(1-e^{-R^{(2)}_{N}/R^{*}_{N}})\\ &\leftrightarrow\frac{N}{U^{(2)}_{N}}-1\ll R^{*}_{N}(1-e^{-R^{(2)}_{N}/R^{*}_{N}})\\ &\leftrightarrow R^{(2)}_{N}/R^{*}_{N}\ll 1-e^{-R^{(2)}_{N}/R^{*}_{N}}\\ \end{split}$ (8)
But this is impossible: by Bernoulli’s inequality, $1+x\leq e^{x}$ for all $x$. This implies that $x\leq e^{x}-1$, which in turn implies $-x\leq 1-e^{x}$, which then implies $x>1-e^{-x}$. As it therefore cannot be the case that $R^{(2)}_{N}/R^{*}_{N}<1-e^{-R^{(2)}_{N}/R^{*}_{N}}$, we know that the model cannot be underparameterized if $B/\hat{D}^{\beta}_{1}\ll B/\hat{D}^{\beta}_{2}$.
Now assume that the model is either over- or well-parameterized at 100% of the data (i.e. that $N\geq N_{\text{{opt}}}(U_{D})$): then $U^{(1)}_{N}=\text{{min}}\{N,N_{\text{{opt}}}(U_{D})\}=N_{\text{{opt}}}(U_{D})$. It is reasonable to assume that $N_{\text{{opt}}}(U_{D}/2)<N_{\text{{opt}}}(U_{D})$ (i.e. that the model’s optimal number of parameters scales monotonically with the amount of pretraining data), which implies that $U^{(2)}_{N}=\text{{min}}\{N,N_{\text{{opt}}}(U_{D}/2)\}=N_{\text{{opt}}}(U_{D}/2)<N_{\text{{opt}}}(U_{D})=U^{(1)}_{N}$. With a fixed number of parameters $N$, $\hat{N}$ (Equations 5c, 6d-e) is a monotonic function of $U_{N}$ (by construction; Muennighoff et al., 2024), so we have $U^{(1)}_{N}>U^{(2)}_{N}\rightarrow\hat{N}_{1}>\hat{N}_{2}$.
But this implies that $A/\hat{N}^{\alpha}_{1}<A/\hat{N}^{\alpha}_{2}$. This—along with the assumption that $B/\hat{D}^{\beta}_{1}\ll B/\hat{D}^{\beta}_{2}$—contradicts the observed result that $L^{(4)}(N,U_{D})\approx L^{(4)}(N,U_{D}/2)$, and so we know that the model cannot be well- or over-parameterized if $B/\hat{D}^{\beta}_{1}\ll B/\hat{D}^{\beta}_{2}$.
Given that it cannot be the case that the model of the 100% run is under-, well-, or over-parameterized if $B/\hat{D}^{\beta}_{1}\ll B/\hat{D}^{\beta}_{2}$ (and that it must be one of the three), it therefore cannot be the case that $B/\hat{D}^{\beta}_{1}\ll B/\hat{D}^{\beta}_{2}$. As argued above, the coefficients $B$ and $\beta$ must be positive, so it also cannot be the case that $B/\hat{D}^{\beta}_{1}\gg B/\hat{D}^{\beta}_{2}$: we must then conclude that $B/\hat{D}^{\beta}_{1}\approx B/\hat{D}^{\beta}_{2}$. As $\hat{D}_{1}\approx 2\hat{D}_{2}$ (see Equation 7), we know that $B/\hat{D}^{\beta}_{1}\approx(B/\hat{D}^{\beta}_{2})/2^{\beta}$, (and that $12^{\beta}>$ 1, as $\beta$ must be positive). It must therefore be the case that $B/\hat{D}^{\beta}_{1}\approx 0$.
Given that $B/\hat{D}^{\beta}_{1}\approx B/\hat{D}^{\beta}_{2}$, the equality in Equation 6a implies that $A/\hat{N}^{\alpha}_{1}\approx A/\hat{N}^{\alpha}_{2}$, which in turn implies that $\hat{N}_{1}\approx\hat{N}_{2}$. If GFoLDS were over- or well-parameterized at 100% of the data, then the monotonicity of $\hat{N}$ would imply that $\hat{N}_{1}>\hat{N}_{2}$ (as discussed above), contradicting the conclusion that $\hat{N}_{1}\approx\hat{N}_{2}$.
On the other hand, replacing the inequalities in Equation 8 with (approximate) equalities yields $R^{(2)}_{N}/R^{*}_{N}\approx 1-e^{-R^{(2)}_{N}/R^{*}_{N}}$: note that the only value of $x$ for which $x=1-e^{-x}$ is 0. If GFoLDS is underparameterized at the 50% run, then $U^{(2)}_{N}=\text{{min}}\{N,N_{\text{{opt}}}(U_{D}/2)\}=N$, which implies that $R^{(2)}_{N}=0$, which in turn implies that $R^{(2)}_{N}/R^{*}_{N}=0$. Given that $N_{\text{{opt}}}(U_{D})$ is (likely) greater than $N_{\text{{opt}}}(U_{D}/2)$, it is also the case that $R^{(1)}_{N}/R^{*}_{N}=0$ and $U^{(1)}_{N}=\text{{min}}\{N,N_{\text{{opt}}}(U_{D})\}=N=U^{(2)}_{N}$: this implies that $A/\hat{N}^{\alpha}_{1}=A/\hat{N}^{\alpha}_{2}$ (see Equations 6d and 6e), which is congruent with the conclusion that $\hat{N}_{1}\approx\hat{N}_{2}$.